OpenStudy (anonymous):

z=2squareroot2 + 2squareroot2i w=2squareroot2 + 2squareroot3i solve z^8/w^6

OpenStudy (anonymous):

\[\frac{ (2\sqrt{2}+2\sqrt{2i})^8 }{ (2\sqrt{2}+2\sqrt{3i})^6 }\]

OpenStudy (anonymous):

Convert to polar, use DeMoivre's theorem.